Write the first four terms of each sequence whose general term is given.
1) a n= 4n-1
2) a n =3n/n+5

Answer by stanbon(75887)

Write the first four terms of each sequence whose general term is given.
1) a n= 4n-1
a1 = 4-1 = 3
a2 = 8-1 = 7
a3 = 12-1 = 11
a4 = 1601 = 15
-----------------------

2) a n =3n/n+5
a1 = 3/6 = 1/2
a2 = 6/7
a3 = 9/8
a4 = 12/9
------------
Cheers,
Stan H.